OpenStudy (michellemn):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (michellemn):

@iGreen

OpenStudy (sleepyjess):

First, what would 40% of .50 be?

OpenStudy (michellemn):

0.20

OpenStudy (sleepyjess):

Yes, so the sale price would be .70 correct?

OpenStudy (michellemn):

Yes!

OpenStudy (sleepyjess):

But the profit on that would only be .2. So we need to use the equation 500=.2x where x is the number of flowers.

OpenStudy (michellemn):

So, The answer would be 2500?

OpenStudy (sleepyjess):

Yes, 2500=flowers

OpenStudy (anonymous):

o they sell a flower at 140% from the original price that would be 0.50$......100% x$.............140% x= 0.50*140/100=0.70$ a flower to make 500$ you shoul sell 500/0.70= 714 flower

OpenStudy (sleepyjess):

No, it has to be $500 in profit.

OpenStudy (sleepyjess):

Not just $500.

OpenStudy (anonymous):

p= = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower Then P = mcn - cn Rearrange to find n P = n(mc - c) n = P / (mc - c) Substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 They must sell 2500 flowers to make a profit of $500.

OpenStudy (michellemn):

But the question says "They set the purchase price of each flower at the purchase price PLUS 40%." So should it be 0.50 plus 40% so It would be $0.70 a flower not $0.20??????

OpenStudy (michellemn):

@sleepyjess

OpenStudy (sleepyjess):

But they only make .20 profit on each flower